OpenStudy (anonymous):

Please help........... Suppose we start with unlimited supplies of a 20% h2so4 solution and of a 50% h2so4 solution. We combine x liters of the first with y liters of the second to produce 7 liters of a 40% h2so4 solution. What does x equal?

OpenStudy (welshfella):

work in masses of pure sulphuric acid the mass of acid in 40% solution is 7 * 10 * 40 = 2800g

OpenStudy (anonymous):

2.8?

OpenStudy (welshfella):

x liters of 20% contains 200x g y liters of 50% = 500y g

OpenStudy (welshfella):

so 200x + 500y = 2800

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

whats next

OpenStudy (welshfella):

now we need to find y in terms of x. Note that the 50% solution i s 5/2 = 2.5 times as strong as the 20% So y liters of 50% = how many liters of 20%?

OpenStudy (anonymous):

where did u get the 5/2?

OpenStudy (welshfella):

the ratio of 50/20 = 2/2

OpenStudy (welshfella):

* 5/2

OpenStudy (anonymous):

oh okay

OpenStudy (welshfella):

so y liters of 50% = 2.5x liters of 20%

OpenStudy (anonymous):

yea

OpenStudy (welshfella):

200x + 500*2.5x = 2800 solve for x

OpenStudy (anonymous):

1.93?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

hmmmm it's supposed to be 2.33333333

OpenStudy (welshfella):

oh yes I've gone wrong somewhere

OpenStudy (welshfella):

something wrong in my logic somewhere

OpenStudy (dumbcow):

set up 2 equations... x + y = 7 20x + 50y = (40)(7)

OpenStudy (welshfella):

yep thats the way

OpenStudy (anonymous):

hmmm and then plug it in huh? y=7-x

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

got it thanx

OpenStudy (welshfella):

yea thats an easier way to do it. I still can't see what i did wrong . Worrying !

OpenStudy (welshfella):

i see now!

OpenStudy (welshfella):

200x + 500y = 2800 was correct then next step was to replace y with 7 - x

OpenStudy (welshfella):

- that is y in tems of x.

OpenStudy (welshfella):

I had a mental aberration!